OpenStudy (anonymous):

If tan(x) = 3/2, what does sec(x) = in quadrant 3?

OpenStudy (mathmale):

Recall that the tangent function is defined as \[\tan \Theta=\frac{ opp }{ adj }\] and the secant as\[\sec \theta=\frac{ hyp }{ adj }\]

OpenStudy (anonymous):